# Testing the framework of the halo occupation distribution with assembly bias modeling

### v1

@@ Abstract
pubyear: 2020pagerange: Testing the framework of the halo occupation distribution with assembly bias modelingReferences
@@ 1 Introduction
At small scales, the clustering amplitude is also sensitive to the physics underlying galaxy formation and evolution (see, e.g., Zheng et al. 2007; White et al. 2007; Brown et al. 2008 and references therein). Due to the dual dependence on cosmology and galaxy formation it is challenging, but not impossible, to extract cosmological information on small scales. One way to do this is to use N-body simulations to provide an accurate description of the non-linear evolution of the dark matter field. Approximate statistical methods can then be used to place galaxies within dark matter halos, which allows sufficient sampling in the parameter space to marginalize over the model dependence. This combined approach can allow us to extract cosmological information at non-linear scales, but requires N-body simulations spanning the range of cosmologies to be tested. In order to interpolate between a small number of such simulations, the emulator method (Heitmann et al. 2009) has been demonstrated to provide accurate measurements for the growth rate of structure using the recent observational data (see Kwan et al. 2015; Zhai et al. 2019; Chapman et al. 2021; Zhai et al. 2023b; Yuan et al. 2022; Kwan et al. 2023b and references therein), as well as investigations into other aspects (Heitmann et al. 2010; Wibking et al. 2017; Euclid Collaboration et al. 2019; Rogers et al. 2019; McClintock et al. 2019; Nishimichi et al. 2019; Lange et al. 2021; Storey-Fisher et al. 2022; Chapman et al. 2023).
@@
In addition to the standard two point correlation function other summary statistics, providing extra small-scale non-linear information, can be explored with a similar emulator approach. These include the marked correlation function (Storey-Fisher et al. 2022), void-based statistics (Fraser et al., 2024), the density split method for galaxy clustering (Paillas et al. 2024), the wavelet scattering transform algorithm (Valogiannis & Dvorkin 2022; Valogiannis et al. 2024), and the $k-$th nearest neighbor statistics (Yuan et al. 2024). When analysing the two point functions and these statistics, marginalising over the HOD parameters to remove the effects of galaxy formation relies on the HOD being a complete description of the link between galaxies and mass in the Universe. This motivates us to consider the impact of more complications in the model of the galaxy halo connection. This will also allow us to investigate the correlation between the empirical models of galaxy bias and the cosmological parameters, and see if any of the extensions we consider can mitigate the tension between the large-scale structure (LSS) measurement and CMB observations. We focus on the parameterization and modeling for both the central and satellite galaxies, including the dark matter halo shape and distribution of central galaxies within the halos. We also test the assumption of the completeness of the HOD model by incorporating an additional parameter to describe model incompleteness.
@@ 3.1.2 Modeling of halo shape in HOD framework
To see the impact of this model, we apply this parameterization to the test samples of the Aemulus suite with large volumes and present the results of $\xi_{0}$ for 100 independent HOD models with various cosmological simulations in Figure 4. The result is represented as the ratio of $\xi_{0}$ with the new satellite distribution divided by the measurement of the original HOD mock. We can see that these additional degrees of freedom can make the clustering amplitude lower than assuming a spherical distribution of satellites. The impact is clearly scale-dependent and more significant in the one-halo regime as expected. The overall result shows that the ellipsoidal halos are less clustered than spherical halos. This result is not quite consistent with the previous analysis based on the MDPL2 simulation shown in Figure 2, which shows that the sphericalization can reduce the clustering amplitude for the same statistics. This is likely caused by the artifact of the rotation operation since it can increase the separation of satellites within the dark matter halos on average as pointed out in van Daalen et al. (2012). Nevertheless, our model is close to the findings of Faltenbacher & White (2010), which shows that the near-spherical halos are more clustered than non-spherical ones but this study is in the framework of assembly bias instead of empirical parameterization, by simply splitting and comparing galaxies based on sphericity and other properties. Compared with those extensive discussions of the dependency on halo mass, concentration, and halo spin, our parameterization only serves as a simplified approximation.
@@ 3.3 Baryonic effects
In the left-hand panel of Figure 9, we show the constraint on the growth rate related parameters considering the baryonic effect and compare them with the reference results. It shows that the cosmological inference is almost unchanged, indicating that the approximate model doesn’t affect the measurement of $f\sigma_{8}$. On the other hand, the constraint on the parameter $B_{f}$ is centered at unity. We do not see a noticeable baryonic effect from this clustering analysis. In order to see if this result is due to a statistical fluke or not, we also run tests with the new emulator on the CMASS galaxies split by different properties and find that although the growth rate measurement is not significantly affected (Figure 7), the constraint on $B_{f}$ is somewhat higher than Aemulus V. We display the constraint in the left panel of Figure 10. This seems to show that the halo mass prediction from the N-body simulation is underestimated. However, the deviation from unity is less than $2\sigma$, and thus the finding is not conclusive.
On the other hand, we note that our scaling is only on the halo mass without considering effects on the internal structures of the dark matter halos such as the radius or radial density profile. This may induce potential self-inconsistency or artifacts and may explain the constraint on $B_{f}$ since the cosmological measurement from clustering is also from the velocity field, which is not modified in our model. Additionally, the change in dark matter halo mass behaves exactly as the the effect of assembly bias, but our model is more simplified and can absorb any dependency on parameters rather than halo mass. A more realistic approach could mimic models like the decorated HOD that can redistribute the halo catalog during mock construction as presented in Hearin et al. (2016).
@@
Figure 9: Similar to Figure 6 but for baryonic model (left) and position bias of central galaxies (right). The result from Aemulus 5 and Aemulus $\nu$ is also shown for reference.
Figure 10: Constraint on the extended model parameter. Left: baryonic parameter $B_{f}$; Right: position bias of central galaxies. The result shows the measurements from the BOSS galaxies split by different properties.
@@ Acknowledgements
ZZ appreciates helpful comments and discussions with Renyue Cen, Xiaodong Li, Yi Zheng, Jeremy Tinker and the Aemulus team. ZZ is supported by NSFC (12373003), the National Key&D Program of China (2023YFA1605600), and acknowledges the generous sponsorship from Yangyang Development Fund.

### v2

@@ Abstract
pubyear: 2020pagerange: Testing the framework of the halo occupation distribution with assembly bias modeling and empirical extensionsReferences
@@ 1 Introduction
At small scales, the clustering amplitude is also sensitive to the physics underlying galaxy formation and evolution (see, e.g., Zheng et al. 2007; White et al. 2007; Brown et al. 2008; Wechsler & Tinker 2018; Salucci 2019 and references therein). Due to the dual dependence on cosmology and galaxy formation it is challenging, but not impossible, to extract cosmological information on small scales. One way to do this is to use N-body simulations to provide an accurate description of the non-linear evolution of the dark matter field. Approximate statistical methods can then be used to place galaxies within dark matter halos, which allows sufficient sampling in the parameter space to marginalize over the model dependence. This combined approach can allow us to extract cosmological information at non-linear scales, but requires N-body simulations spanning the range of cosmologies to be tested. In order to interpolate between a small number of such simulations, the emulator method (Heitmann et al. 2009) has been demonstrated to provide accurate measurements for the growth rate of structure using the recent observational data (see Kwan et al. 2015; Zhai et al. 2019; Chapman et al. 2021; Zhai et al. 2023b; Yuan et al. 2022; Kwan et al. 2023b and references therein), as well as investigations into other aspects (Heitmann et al. 2010; Wibking et al. 2017; Euclid Collaboration et al. 2019; Rogers et al. 2019; McClintock et al. 2019; Nishimichi et al. 2019; Lange et al. 2021; Storey-Fisher et al. 2022; Chapman et al. 2023).
@@
In addition to the standard two point correlation function other summary statistics, providing extra small-scale non-linear information, can be explored with a similar emulator approach. These include the marked correlation function (Storey-Fisher et al. 2022), void-based statistics (Fraser et al., 2024), the density split method for galaxy clustering (Paillas et al. 2024), the wavelet scattering transform algorithm (Valogiannis & Dvorkin 2022; Valogiannis et al. 2024), and the $k-$th nearest neighbor statistics (Yuan et al. 2024). When analysing the two point functions and these statistics, marginalising over the HOD parameters to remove the effects of galaxy formation relies on the HOD being a complete description of the link between galaxies and mass in the Universe. This motivates us to consider the impact of more complications in the model of the galaxy halo connection. This will also allow us to investigate the correlation between the empirical models of galaxy bias and the cosmological parameters, and see if any of the extensions we consider can mitigate the tension between the large-scale structure (LSS) measurement and CMB observations. We consider the parameterization and modeling of both the central and satellite galaxies, including the dark matter halo shape and distribution of central galaxies within the halos. We also test the assumption of the completeness of the HOD model by incorporating an additional parameter to describe model incompleteness. Although these effects may not directly result from the assembly history of dark matter halos, they are inspired by considering how galaxy formation might be affected by halo properties. From this point of view, any impact on the clustering property that is not fully modeled by halo mass could be understood as a secondary bias or assembly bias.
@@ 3.1.2 Modeling of halo shape in HOD framework
To see the impact of this model, we apply this parameterization to the test samples of the Aemulus suite with large volumes and present the results of $\xi_{0}$ for 100 independent HOD models with various cosmological simulations in Figure 4. The result is represented as the ratio of $\xi_{0}$ with the new satellite distribution divided by the measurement of the original HOD mock. We can see that these additional degrees of freedom can make the clustering amplitude lower than assuming a spherical distribution of satellites. The impact is clearly scale-dependent and more significant in the one-halo regime as expected. The overall result shows that the ellipsoidal halos are less clustered than spherical halos. This result is related to the work of Faltenbacher & White (2010), which showed that near-spherical halos are more clustered than non-spherical ones. Their study was in the framework of assembly bias rather than empirical parameterization: they simply split the galaxy population and comparing their clustering amplitude. This result should not be confused with the process of sphericalization, as shown in Figure 2. Sphericalization can reduce the clustering amplitude for the same statistics ($\xi_{0}$). This is likely caused by the rotation operation increasing the separation of satellites within the dark matter halos on average (van Daalen et al., 2012). Our model works in a different way with the parameterization in Eq. (7) and (3.1.2) preserving the volume of the spheroid and ellipsoid, and thus the average separation of the galaxy pairs of the one-halo terms. The resultant behavior of the clustering statistics becomes closer to the physical exploration in Faltenbacher & White (2010). We note that compared with those extensive discussions of the dependency on halo mass, concentration, and halo spin, our parameterization only serves as a simplified approximation.
@@ 3.3 Baryonic effects
In the left-hand panel of Figure 9, we show the constraint on the growth rate related parameters considering the baryonic effect and compare them with the reference results. It shows that the cosmological inference is almost unchanged, indicating that the approximate model doesn’t affect the measurement of $f\sigma_{8}$. On the other hand, the constraint on the parameter $B_{f}$ is centered at unity. We do not see a noticeable baryonic effect from this clustering analysis. In order to see if this result is due to a statistical fluke or not, we also run tests with the new emulator on the CMASS galaxies split by different properties and find that although the growth rate measurement is not significantly affected (Figure 7), the constraint on $B_{f}$ is somewhat higher than Aemulus V. We display the constraint in the left panel of Figure 10. Because $B_{f}$ is a scaling factor for halo mass, the preference for $B_{f}>1$ indicates that the halo mass prediction from the N-body simulation is somewhat underestimated. However, the deviation from unity is less than $2\sigma$, and thus the finding is not conclusive. In addition, $B_{f}$ is correlated with other parameters, affecting both centrals and satellites. The high value of $B_{f}$ correlates with slightly lower $f_{max}$, the incompleteness parameter of central occupancy at the massive end, and lower $\eta_{vs}$, the velocity bias of satellites and slightly higher $\alpha$ for satellite occupancy. In general, a high $B_{f}$ can produce a galaxy mock with higher satellite fraction and thus lower number of centrals in order to preserve the galaxy number density. To maintain the galaxy bias anchored by a large scale measurement, the lower number of centrals has to be offset by less occupancy of the more massive halos and thus a lower value of $f_{max}$. Since the overall impact of $B_{f}$ is relatively weak at large scales where the centrals dominate (Fig. 8), this influence is not very significant. For the same reason, a high value of $B_{f}$ can limit satellites to a narrower range of halo mass (a halo before scaling without satellites can host satellites after scaling, and there are fewer halos available at the massive end due to $f_{max}$). The model has to produce more satellite pairs with a faster occupancy (higher $\alpha$) as a function of halo mass. In order to balance the enhanced clustering due to more one-halo terms, the velocity field needs to be suppressed to fit the data.
This interpretation does not consider effects on the internal structures of the dark matter halos such as the radius or radial density profile. These may partially explain the constraint on $B_{f}$ since the cosmological measurement from clustering also depends on the velocity field, which is not modified in our model. Additionally, the change in dark matter halo mass behaves exactly as that due to assembly bias, but our model is more simplified and can absorb any dependency on parameters rather than halo mass. A more realistic approach could mimic models like the decorated HOD that can redistribute the halo catalog during mock construction as presented in Hearin et al. (2016).
@@
Figure 9: Similar to Figure 6 but for baryonic model (left) and position bias of central galaxies (right). The results from Aemulus 5 and Aemulus $\nu$ are also shown for reference.
Figure 10: Constraint on the extended model parameter from the BOSS galaxies split by different properties. Left: baryonic parameter $B_{f}$; Right: position bias of central galaxies.
@@ Acknowledgements
We thank the anonymous reviewers for the helpful and insightful comments and suggestions that have significantly improved this paper. ZZ appreciates helpful comments and discussions with Renyue Cen, Xiaodong Li, Yi Zheng, Jeremy Tinker and the Aemulus team. ZZ is supported by NSFC (12373003), the National Key&D Program of China (2023YFA1605600), and acknowledges the generous sponsorship from Yangyang Development Fund.